OpenStudy (anonymous):

What is the solution to the equation? 6x – 27 = –3 A. x = –5 B. x = –4 C. x = 4 D. x = 5

OpenStudy (anonymous):

6x - 27 = -3 + 27 + 27 6x = 24 24 / 6 = x x = 4

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

Yep :)